Theme Year in Mathematics, 2003-2004 Inverse Problems and Related Topics

Helsinki University of Technology Institute of Mathematics

January 13, 2003

Conferences

• 18.–22.8.2003: Jyväskylä Summer School on inverse problems.

The international 13. Jyväskylä Summerschool has inverse problems as the mathematics topics. There are two courses, given by Giovanni Alessandrini (Trieste) and Liliana Borcea (Rice).

• 25.–29.8.2003: Workshop on Geometry and Analysis in Inverse

Problems The participants will include Carlos Alves, Elena Beretta, David Colton, Allan Greenleaf, Martijn de Hoop, Elisa Francini, Alberto Grunbaum, Alexander Katchalov, Yaroslav V. Kurylev, Rainer Kress, Cliff Nolan and Gunther Uhlmann.

• 31.5.–4.6.2004: Theory and Applications in Inverse Problems

International Organizing Committee

• David Colton (Delaware)

• Rainer Kress (Göttingen)

• Yaroslav V. Kurylev (Loughborough)

• Gunther Uhlmann (Washington)

Organizing Committee

• Jari Kaipio

• Matti Lassas

• Petri Ola

• Lassi Päivärinta

• Erkki Somersalo